a) If a single six-sided die is rolled five times, what is the probability that a 6 is thrown exactly three times?

b) A person receives an average of one e-mail message per half-hour interval. Assume that e-mails are received randomly in time, find the probabilities that in a particular hour 0,1,2,3,4,5 messages are received.
Answer #1

a) P(Getting 6 in any single throw) = 1/6

Hence,

P(Getting 6 three times in 5 throws)

2

= 0.03215

b) Average = 1 email per half hour = 2 emails per hour

Hence,

x Probability calculation Probability
0 \frac{e^{-2}2^0}{0!} 0.1353
1 \frac{e^{-2}2^1}{1!} 0.2707
2 \frac{e^{-2}2^2}{2!} 0.2707
3 \frac{e^{-2}2^3}{3!} 0.1804
4 \frac{e^{-2}2^4}{4!} 0.0902
5 \frac{e^{-2}2^5}{5!} 0.0361